OpenStudy (anonymous):

Solve quadratic equation by completing the square. Check both roots. problem below

OpenStudy (anonymous):

\[2x^2-12x +16=0\]

OpenStudy (anonymous):

please explain each step I am having trouble with it

OpenStudy (anonymous):

2(x^2-6x+8)=0 2[(x-4)(x-2)]=0

OpenStudy (anonymous):

please do explain why you do each step

OpenStudy (nottim):

1st step- Factor out 2. Second step- What 2 numbers add to the end value, and multiply to the middle value?

OpenStudy (nottim):

Those 2 values are then put into the brackets as you see them.

OpenStudy (nottim):

And here, the signs of positive and negatives are reversed because when you solve for x individually, x-4=0 x=4. Ok?

OpenStudy (anonymous):

ok but there is supposed to be two answers

OpenStudy (anonymous):

4 is one of them

OpenStudy (anonymous):